Antiques R Us is a mature manufacturing firm. The company just paid a dividend of $10.70, but management expects to reduce
Answer #1

Current price=D1/(Required return-Growth rate)

=(10.7*(1-0.05))/(0.12-(0.05)

=(10.165/0.17)

=$59.79(Approx).